Please help. A backyard has dimentions 5 1/4 yards by 4 2/3 yards. What is the area of the backyard in square yards? I am having trouble finding the answer.
Answer by checkley71(8403)

THE AREA OF A RECTANGLE (BACK YARD)IS
AREA=LENGTH*WIDTH
AREA=(5 1/4)(4 2/3) ITS VERY DIFFICULT TO MANIPULATE MIXED FRACTION THUS WE
CONVERT THEM TO IMPROPER FRACTION. THUS 5 1/4=21/4 & 4 2/3=14/3.
NOW WE CAN SET UP THE EQUATION & MULTIPLY THE FRACTIONS.
AREA=21/4*14/3
NOW WE MULTIPLY NUMERATORS AND DENOMINATORS TO GET
AREA=294/12
AREA=24.5 SQUARE YARDS IS THE SIZE OF THIS YARD.